OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (goformit100):

You can Smartly use Geometry to solve Questions like This. Now Try to use it, and let me Know if you are done.

OpenStudy (anonymous):

Im not in geometry! im in algebra still!

OpenStudy (aripotta):

what is point-slope form

OpenStudy (anonymous):

\[y - y _{1} = m(x - x _{1})\]

OpenStudy (aripotta):

ok, so just substitute the x and y from (3, -2) for y1 and x1, and the slope, 2, for m

OpenStudy (anonymous):

y - (-2) = 2(x - 3)

OpenStudy (aripotta):

so y + 2 = 2(x - 3)

OpenStudy (aripotta):

do you know how to change it into slope-intercept form?

OpenStudy (anonymous):

y + 2 = 2(x - 3) y + 2 = 2x - 6 - 2 - 2 y = 2x - 8

OpenStudy (anonymous):

is this right?

OpenStudy (aripotta):

yes :D

OpenStudy (anonymous):

i just dont kno bout standard form

OpenStudy (aripotta):

do you know what standard form is?

OpenStudy (anonymous):

no...

OpenStudy (aripotta):

Ax + By = C

OpenStudy (aripotta):

and A can't be negative or a decimal

OpenStudy (anonymous):

i dont kno how to solve that :(

OpenStudy (aripotta):

y = 2x - 8 you have to subtract 2x -2x + y = -8 but A can't be negative, so you then have to divide everything by -1 2x - y = 8

OpenStudy (anonymous):

equation is y=2x-8

OpenStudy (anonymous):

so you divide 2x and 8 by -1?

OpenStudy (anonymous):

point slope equation is y+2 = 2 (x-3)

OpenStudy (aripotta):

you'd have to divide the whole equation by -1. and all that does is change the signs

OpenStudy (anonymous):

y = 2 x-8 that is slope intercept form

OpenStudy (anonymous):

standard form is y+2 = 2 (x-3)

OpenStudy (anonymous):

so its -2x - y = -8

OpenStudy (aripotta):

no it's 2x - y = 8

OpenStudy (anonymous):

I just gave you all three forms that you need

OpenStudy (anonymous):

ohhhh okay thank you :)

OpenStudy (aripotta):

no problemo :]